show that if a rod held at angle Theta to the horizontal and released ,its lower end will not slip if the friction coefficient between rod and ground is greater then 3sinthetacostheta/1+3sin^2theta.
1
Expert's answer
2020-09-06T17:22:40-0400

"\alpha= \frac{mg\frac{1}{2}cos\theta}{ \frac{ml^2}{3}}= \frac {3 g\cos\theta}{2\times3}"


"a_C= \frac {1}{2}\alpha= \frac {3}{4}g \cos\theta"

Now "\mu N=ma_C \sin\theta or \mu N= \frac {3}{4} mg \sin\theta \cos\theta..........(i)"


Further, "mg-N=ma_y" or "N=mg- ma_y \cos \theta" or "N= mg- \frac {3}{4} mg \cos^2 \theta..........(ii)"


Dividing equation (i) and (ii) we have


"\mu= \frac {\frac{3}{4}\cos\theta\sin\theta} {1-\frac{3}{4}\cos^2\theta}" = "\frac {3\sin\theta \cos \theta}{4-3 \cos^2\theta}"


= "\frac {3\sin\theta \cos\theta}{1+3\sin^2\theta}" Hence Proved.